40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 1 16:06:32 2014, Koji, Configuration, LSC, DRMI locked 6x
    Reply  Sat Nov 8 16:01:32 2014, ericq, Configuration, LSC, DRMI sensing 6x
       Reply  Mon Nov 10 18:23:10 2014, ericq, Configuration, LSC, DRMI sensing REFL11.pngREFL33.pngREFL55.pngREFL165.pngAS55.png
Message ID: 10690     Entry time: Sat Nov 8 16:01:32 2014     In reply to: 10661     Reply to this: 10693
Author: ericq 
Type: Configuration 
Category: LSC 
Subject: DRMI sensing 

Here are some preliminary results from the sensing sweeps I did the other night. 

Notes:

  • The analog AS55I signal chain is almost certainly busted in some way. This would also explain the odd looking error signals in SRX, and was actually hypothesized by Koji when discussing the SRX oddness. 
  • I used the same mirror calibration numbers from Koji's recent Elogs to turn these into counts/m.
  • MICH was excited via differential ITM motion.  I also performed a TF with BS driven MICH, with the compensating PRM output matrix in place, and it looks different, but I haven't looked too deeply into it yet. 
  • The angles plotted are in regard to the analog I and Q signals (i.e., I took TFs to I_ERR and Q_ERR and then unrotated by the digital rotation angle); this is why I suspect AS55I is broken, as all of the signals are entirely in the analog Q.
  • The amplitudes seem to be roughly consistent with Koji's recent observations. 
  • I still need to cut out the violin-filter-corrupted data points to quote the sensing elements with error bars...

Plots!

 REFL11.pngREFL33.png

REFL55.pngREFL165.png

AS55.png

xml files, and DttData matlab script used to generate these plots is attached. 

Attachment 6: DRMIsensing.zip  362 kB
ELOG V3.1.3-